Circular No. 8455

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2004 V6–V12 (SOHO)

Additional SOHO comets (continuation to IAUC 8451):

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 V6 Nov. 4.113 14^{h}29.^{m}5 −17^{o}16′ C2  TH 2004-X72
C/2004 V7  4.146 14 33.8 −16 43 C3/2 TH 2004-X72
C/2004 V8  7.696 14 39.3 −19 12 C3/2 KB 2004-X72
C/2004 V9  7.754 14 55.8 −18 32 C2/3 HO 2004-X73
C/2004 V10  8.129 14 49.7 −16 43 C2  KB 2004-X73
C/2004 V11  8.746 14 53.3 −18 12 C2  TH 2004-X74
C/2004 V12  10.029 14 40.6 −21 18 C3/2 XL 2004-X74

COMET C/2004 V13 (SWAN) On Nov. 30, M. Mattiazzo (Adelaide, S. Australia) reported that he

noticed images of a faint object moving on SWAN images (his positions below have uncertainties of a degree or more due to the poor resolution of the ultraviolet imager on SWAN/SOHO), speculating then that, if real, the object might become visible in the days ahead in the C3 coronagraph. Today, S. Hönig has reported the appearance of a comet with a tail in C3 images, which K. Battams reports has brightened slightly from mag ∼ 6.5 on Dec. 16.26 UT to ∼ 6.1 on Dec. 16.70–16.74. While only the first SOHO position by Battams is given below (reduced by B. G. Marsden), a span of 14 hours of SOHO astrometry appears on MPEC 2004-Y02.

2004 UT  α_{2000}  δ_{2000}  Observer
Nov. 9  12^{h}48^{m}  −19.^{o}5  SWAN
16  13 15  −20.5  _{”}
21  13 41  −21.0  _{”}
25  14 08  −21.5  _{”}
Dec. 16.154 17 04.9 −19 54′  SOHO

It does appear that the SWAN object is identical with the SOHO object, and the following very preliminary parabolic orbital elements by Marsden also appear on MPEC 2004-Y02.

T = 2004 Dec. 21.139 TT  ω = 94.^{o}253
Ω = 207.015 }2000.0
q = 0.17727 AU  i = 34.463
2004 December 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green